# Inductive Gradient Adjustment for Spectral bias in implicit neural representations

### v1

@@ 3.2 Improving spectral bias via Inductive Gradient Adjustment
Although $\tilde{{\bm{K}}}$ avoids intractability, it still encounters the curse of dimensionality as $N$ increases. In Theorem 4.2, we have that training dynamics of ${\bm{r}}_{t}$ could be estimated via inductive generalization of the linear dynamics model of sampled data points and analyzable error. Inspired by this theoretical foundation, we propose a practical inductive gradient adjustment (IGA) method. Specifically, assumed that $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ have been sorted based on the proximity of input samples, we divide these samples into $n$ groups, where each group is ${\bm{X}}^{j}=\{({\bm{x}}^{j}_{i},y^{j}_{i})\}_{i=1}^{p}$. Then we sample one point from each group to form the sample set ${\bm{X}}_{e}$. We denote the eNTK matrix on ${\bm{X}}_{e}$ as the empirical inductive NTK matrix $\tilde{{\bm{K}}}_{e}$ and construct the corresponding transformation matrix $\tilde{{\bm{S}}}_{e}$. Then inductive adjustments are generalized to gradients of whole data points as follows:
$\Theta_{t+1}=\Theta_{t}-\eta\sum_{i=1}^{p}\nabla_{\Theta_{t}}f({\bm{X}}_{i},\Theta_{t})\tilde{{\bm{S}}}_{e}{\bm{r}}^{i}_{t},$ (4)
@@
where ${\bm{X}}_{i}=\{({\bm{x}}^{j}_{i},y^{j}_{i})\}_{j=1}^{n}$ and ${\bm{r}}^{i}_{t}$ denotes the corresponding residual vector at time step $t$; $\tilde{{\bm{S}}}_{e}=\sum_{i=1}^{N}(g_{i}(\tilde{\tilde{\lambda}}_{i})/\tilde{\tilde{\lambda}}_{i})\tilde{\tilde{{\bm{v}}_{i}}}\tilde{\tilde{{\bm{v}}_{i}}}^{\top}$; $\tilde{{\bm{K}}}_{e}=\nabla_{\Theta_{t}}f({\bm{X}}_{e};\Theta_{t})^{\top}\nabla_{\Theta_{t}}f({\bm{X}}_{e};\Theta_{t})=\sum_{i=1}^{N}\tilde{\tilde{\lambda}}_{i}\tilde{\tilde{{\bm{v}}_{i}}}\tilde{\tilde{{\bm{v}}_{i}}}^{\top}$. As shown in equation 4, all ${\bm{r}}^{i}_{t}$ are linearly transformed by $\tilde{{\bm{S}}}_{e}$. Please note that for each vector ${\bm{r}}^{i}_{t}$ (where $i=1,\dots,p$), its $j$-th element is scaled by the $j$-th column of the matrix $\tilde{{\bm{S}}}_{e}$. And the $j$-th element of these vectors ${\bm{r}}^{i}_{t}$ all belong to the group ${\bm{X}}_{j}$. Therefore, our inductive gradient adjustments are generalized to other data points in each group ${\bm{X}}^{j}$. From our Theorem 4.2, although the inductive generalization by $\tilde{{\bm{K}}}$ introduces errors; these errors can be reduced by increasing data similarity, and wider MLPs also provide a chance to reduce these errors. Further, in most INR tasks, our experimental results illustrate that these errors could be neglected.
@@ 3.3 Implementation Details
Construction Strategy. As we have introduced in equation 2, we construct the transformation matrix by the desired spectrum. We have that $\tilde{{\bm{K}}}_{e}=\sum_{i=1}^{n}\lambda_{i}{\bm{v}}_{i}{\bm{v}}_{i}^{\top}$ and assume that $\lambda_{1}>\dots>\lambda_{n}>0$. To purposefully improve the spectral bias, we balance the eigenvalues of different eigenvectors; impacts are tailored by managing the number of balanced eigenvalues. For vanilla gradient descent, we have $\tilde{{\bm{S}}}_{e}$ as follows:
$\tilde{{\bm{S}}_{e}}=\sum_{i=start}^{end}\frac{\lambda_{start}}{\lambda_{i}}\mathbf{v}_{i}\mathbf{v}_{i}^{\top}+\sum_{i\notin[start,end]}1\cdot\mathbf{v}_{i}\mathbf{v}_{i}^{\top}.$ (5)
Generally, $start$ is fixed at $1$; $end$ represents the controlled spectral range. The larger the value of $end$, the stronger impacts over spectral bias.
Due to the introduction of adaptive learning rates and momentum in Adam, training dynamics of MLPs becomes exceptionally challenging and remains an open problem. Intuitively, momentum is the linear combination of previous gradients, which implies that momentum has the similar direction with current adjusted gradients. This inspires us to extends the equation 4 to Adam. Adaptive learning rates typically result in larger update steps for parameters (Kingma, 2014; Wilson et al., 2017; Reddi et al., 2019). Therefore, we utilize $\lambda_{end}$ to balance eigenvalues in equation 5.
Sampling Strategy. As illustrated in Sec. 3.2, we need to sample data points from groups to compute $\tilde{{\bm{K}}}$. Discussing all sampling strategies is beyond the scope of our work. For coordinate inputs, we partition inputs by Euclidean distance of samples into $n$ groups with $p$ points and largest residual points in each group are sampled. Details of hyperparameters $n$ and $p$ for different settings will be introduced in the experimental section. We provide ablation experiments in our Appendix B to analyze the effect of group size $p$ and points.
@@
Multidimensional output approximation. In practical INRs tasks, the output of MLPs is multidimensional in certain scenarios, such as color images. Compared to the dimension of data points, it is trivial but still affects the efficiency. Therefore, we simply compute the Jacobian matrix by sum-of-logits which has been proved efficient in (Mohamadi et al., 2023; Shi et al., 2024b).
4 Theoretical Analysis
In this section, we firstly introduce our analysis framework, referring to previous works (Du et al., 2018; Arora et al., 2019; Lee et al., 2019; Geifman et al., 2023). Building on this, we prove that adjustments based on eNTK matrix asymptotically converge to those based on NTK matrix as the network width increases. Then we delve into the estimate of training dynamics via inductive generalization of linear dynamics model of sampled data points.
Analysis Framework. For better analysis of NTK matrix, we refer previous works (Du et al., 2018; Arora et al., 2019; Lee et al., 2019; Geifman et al., 2023) and perform a theoretical analysis on the basis of the following settings: assuming that the training set $\{{\bm{x}}_{i},y_{i}\}_{i=1}^{N}$ is contained in one compact set, a two-layer network $f({\bm{x}};\Theta)$ with $m$ neurons is formalized as follows to fit these data:
$f({\bm{x}};\Theta)=(1/\sqrt{m})\sum\nolimits_{r=1}^{m}a_{r}\sigma({\bm{w}}^{\top}_{r}{\bm{x}}+b_{r}),$ (6)
where the activation function $\sigma$ satisfies that $|\sigma(0)|,\|\sigma^{\prime}\|_{\infty},sup_{x\neq x^{\prime}}|\sigma^{\prime}(x)-\sigma^{\prime}(x^{\prime})|/|x-x^{\prime}|<\infty$. The detailed initialization scheme can be found in our Appendix A.1. The loss function is $\frac{1}{2}\sum_{i=1}^{N}(f({\bm{x}}_{i},\Theta_{t})-y_{i})^{2}$. We assume that ${\bm{K}},\tilde{{\bm{K}}}$ are full rank, indicating that they are positive definite. This assumption generally holds due to the complexity of neural networks. Then we have the following standard orthogonal spectral decomposition: ${\bm{K}}=\sum_{i=1}^{N}\lambda_{i}{\bm{v}}_{i}{\bm{v}}_{i}^{\top}$ and $\tilde{{\bm{K}}}=\sum_{i=1}^{N}\tilde{\lambda}_{i}\tilde{{\bm{v}}_{i}}\tilde{{\bm{v}}_{i}}^{\top}$, which eigenvalues are indexed in descending order of magnitude. We construct the corresponding transformation matrices as ${\bm{S}}=\sum_{i=1}^{N}(g_{i}(\lambda_{i})/\lambda_{i}){\bm{v}}_{i}{\bm{v}}_{i}^{\top}$ and $\tilde{{\bm{S}}}=\sum_{i=1}^{N}(g_{i}(\tilde{\lambda}_{i})/\tilde{\lambda}_{i})\tilde{{\bm{v}}}_{i}\tilde{{\bm{v}}}_{i}^{\top}$. Although $\lim_{m\rightarrow\infty}\|{\bm{K}}-\tilde{{\bm{K}}}\|_{F}=0$ (Jacot et al., 2018; Lee et al., 2019; Geifman et al., 2023), it is non-trivial to show that ${\bm{S}},\tilde{{\bm{S}}}$ have the similar impacts on the spectrum. Thanks to the well-established theory of matrix perturbation analysis (Yu et al., 2015; Davis & Kahan, 1970; Baumgärtel, 1984), we have the following Theorem 4.1.
Theorem 4.1.
The following standard orthogonal spectral decomposition exists: ${\bm{K}}=\sum_{i=1}^{N}\lambda_{i}{\bm{v}}_{i}{\bm{v}}_{i}^{\top}$ and $\tilde{{\bm{K}}}=\sum_{i=1}^{N}\tilde{\lambda_{i}}\tilde{{\bm{v}}_{i}}\tilde{{\bm{v}}_{i}}^{\top}$, satisfying that ${\bm{v}}_{i}^{\top}\tilde{{\bm{v}}}_{i}>0$ for $i\in[N]$. We denoted $min\{\lambda_{i}-\lambda_{i+1},\lambda_{i+1}-\lambda_{i+2}\}$ as $G$. Let $\{g_{i}(x)\}_{i=1}^{N}$ be a set of Lipschitz continuous functions, with the Supremum of their Lipschitz constants denoted by $k$, with $\eta<min\{((max(g(\lambda))+min(g(\lambda)))^{-1},(max(g(\tilde{\lambda}))+min(g(\tilde{\lambda})))^{-1}\}$, for $\epsilon>0$, there always exists $M>0$, such that $m>M$, for $i\in[N]$, we have that $|g_{i}(\lambda_{i})-g_{i}(\tilde{\lambda_{i}})|<\epsilon_{1},\|{\bm{v}}_{i}-\tilde{{\bm{v}}}_{i}\|<\epsilon_{2}$; furthermore, we have that:
The detailed proof of Theorem 4.1 can be found in our Appendix A.3. Theorem 4.1 shows that eNTK-based gradient adjustment ($\tilde{{\bm{S}}}$) exhibits almost the same level of control as NTK-based adjustment (${\bm{S}}$) across different feature directions as width increases, thereby leading to almost the same training dynamics of ${\bm{r}}_{t+1}$ and $\tilde{{\bm{r}}}_{t+1}$. More fundamentally, eigenvectors of $\tilde{{\bm{K}}}$ converge one-to-one to eigenvectors of ${\bm{K}}$, thus ensuring that $\tilde{{\bm{S}}}$ impact the similar convergence directions as ${\bm{S}}$. In Theorem 4.2, we show that training dynamics could be estimated by inductive generalization of the linear dynamics model of sampled data points.
Theorem 4.2.
${\bm{X}}=\{{\bm{x}}_{i}\}_{i=1}^{N}$ are partitioned by order into $n$ groups, where each group is ${\bm{X}}_{j}=\{{\bm{x}}^{j}_{1},\dots,{\bm{x}}^{j}_{p}\}$ and $N=np$. For each group, one data point is sampled denoted as ${\bm{x}}^{j}$ and these $n$ data points form ${\bm{X}}_{e}=\{{\bm{x}}^{j}\}_{j=1}^{n},n\ll N$. There exists $\epsilon>0$, for $j=1,\dots,n$ and any $1\leq i_{1},i_{2}\leq p$, such that $|(f({\bm{x}}^{j}_{i_{1}},\Theta_{t})-y^{j}_{i_{1}})-(f({\bm{x}}^{j}_{i_{2}},\Theta_{t})-y^{j}_{i_{2}})|,\|\nabla_{\Theta_{t}}f({\bm{x}}^{j}_{i_{1}})-\nabla_{\Theta_{t}}f({\bm{x}}^{j}_{i_{2}})\|<\epsilon$. Then, for any ${\bm{x}}_{i}$, assumed that ${\bm{x}}_{i}$ belongs to ${\bm{X}}_{j_{i}}$, such that:
$|\Delta f({\bm{x}}_{i},\Theta_{t})-\nabla_{\Theta_{t}}f({\bm{x}}^{j_{i}})^{\top}[p\sum_{j=1}^{n}\nabla_{\Theta_{t}}f({\bm{x}}^{j})(f({\bm{x}}^{j})-y^{j})]|<\epsilon(\frac{(\kappa+n^{3/2})R_{0}+\kappa^{2}}{\sqrt{m}})+\frac{\eta\kappa^{3}R_{0}^{2}}{m^{3/2}},$
where $\Delta f({\bm{x}}_{i},\Theta_{t})=({\bm{r}}_{t+1}-{\bm{r}}_{t})_{i}$ denotes the dynamics of $f({\bm{x}}_{i},\Theta)$ at time step $t$; $R_{0},\kappa$ are constants; $\nabla_{\Theta_{t}}f({\bm{x}}^{j_{i}})^{\top}\nabla_{\Theta_{t}}f({\bm{x}}^{j})$ is the entry in ${\tilde{{\bm{K}}}_{e}}(i,j)$; $\nabla_{\Theta_{t}}f({\bm{x}}):=\nabla_{\Theta_{t}}f({\bm{x}},\Theta_{t})$.
The detailed proof can be found in our Appendix A.4. Theorem 4.2 offers us a new perspective that training dynamics can be estimated by inductive generalization of the linear dynamics model of sampled data points. Moreover, this estimate error decreases as the similarity of data increases, and increasing $m$ provides a chance to reduce this error. Inspired by this, we construct the transformation matrix $\tilde{{\bm{S}}}_{e}$ based on $\tilde{{\bm{K}}}_{e}$ to purposefully improve the spectral bias. Practically, our empirical results show that this error is trivial to purposefully improve the spectral bias.

### v2

@@ 3.2 Inductive Gradient Adjustment
Although adjustment based on $\tilde{{\bm{K}}}$ instead of ${\bm{K}}$ achieves similar effects to ${\bm{K}}$ while avoiding intractability, it still encounters the curse of dimensionality as the data size $N$ increases. One common approach is to estimate $\tilde{{\bm{K}}}$ using sampled data, which is efficient for tasks where $\tilde{{\bm{K}}}$ indirectly impacts training without affecting gradients, such as neural architecture search (Park et al., 2020) or adaptive module (Shi et al., 2024b). However, how to effectively utilize $\tilde{{\bm{K}}}$ from sampled data to directly adjust the training dynamics on the population data remains a problem. In Theorem 3.2, we show that training dynamics of ${\bm{r}}_{t}$ could be estimated via inductive generalization of the linear dynamics model of sampled data and analyzable error. Inspired by this theoretical foundation, we propose to inductively generalize the gradient adjustment based on sampled data to the gradients of the population data. Specifically, considering an input set ${\bm{X}}=\{{\bm{x}}_{i}\}_{i=1}^{N}$ of $N$ coordinates in $\mathbb{R}^{d}$ as the population data, ${\bm{X}}$ can be divided into $n$ groups $(N=np)$, where the $j$-th group ${\bm{X}}^{j}$ and corresponding labels are $\{({{\bm{x}}^{j}_{i}},y^{j}_{i})\}_{i=1}^{p}$. Then we sample one data point from each group to form the sampled set ${\bm{X}}_{e}$, where $|{\bm{X}}_{e}|=n\ll N$. The eNTK matrix on ${\bm{X}}_{e}$ is defined by $\tilde{{\bm{K}}}_{e}=\nabla_{\Theta_{t}}f({\bm{X}}_{e};\Theta_{t})^{\top}\nabla_{\Theta_{t}}f({\bm{X}}_{e};\Theta_{t})$. We construct the corresponding transformation matrix $\tilde{{\bm{S}}}_{e}$. Then inductive adjustments $\tilde{{\bm{S}}}_{e}$ are generalized to gradients of the population data ${\bm{X}}$ as follows:
$\Theta_{t+1}=\Theta_{t}-\eta\sum\nolimits_{i=1}^{p}\nabla_{\Theta_{t}}f({\bm{X}}_{i},\Theta_{t})\tilde{{\bm{S}}}_{e}{\bm{r}}^{i}_{t},$ (4)
where ${\bm{r}}^{i}_{t}$ is the residual of ${\bm{X}}_{i}=\{({{\bm{x}}^{j}_{i}},y^{j}_{i})\}_{j=1}^{n}$; $\tilde{{\bm{S}}}_{e}=\sum_{i=1}^{n}(g_{i}(\tilde{\tilde{\lambda}}_{i})/\tilde{\tilde{\lambda}}_{i})\tilde{\tilde{{\bm{v}}_{i}}}\tilde{\tilde{{\bm{v}}_{i}}}^{\top}$; $\tilde{{\bm{K}}}_{e}=\sum_{i=1}^{n}\tilde{\tilde{\lambda}}_{i}\tilde{\tilde{{\bm{v}}_{i}}}\tilde{\tilde{{\bm{v}}_{i}}}^{\top}$. We introduce the implementation details of Eq. 4 in Sec. 3.4 and summarize the IGA method in the form of an algorithm block 1.
Algorithm 1 Inductive Gradient Adjustment (IGA)
Input: Mini-batch $X_{t}=\{(x_{i},y_{i})\}_{i=1}^{N}$; model $f(\cdot,\Theta_{t})$
Output: Updated parameters $\Theta_{t+1}$
// Sample a subset
Sample $X_{e}\subset X_{t}$ based on the strategy in Sec. 3.4
// Compute gradients and empirical kernel
Compute $G_{e}=\nabla_{\Theta_{t}}f(X_{e};\Theta_{t})$ and then $\tilde{K}_{e}=G_{e}^{\top}G_{e}$
// Construct the transformation matrix
Compute $\tilde{S}_{e}$ based on $\tilde{K}_{e}$ using the method in Sec. 3.4
// Generalize the adjustment to the full mini-batch
Adjusted gradient: $\tilde{g}_{t}=\sum\nolimits_{i=1}^{p}\nabla_{\Theta_{t}}f({\bm{X}}_{i},\Theta_{t})\tilde{{\bm{S}}}_{e}{\bm{r}}^{i}_{t}$; update the parameters by: $\Theta_{t+1}\leftarrow\Theta_{t}-\eta\cdot\tilde{g}_{t}$
return $\Theta_{t+1}$
@@
As shown in equation 4, all ${\bm{r}}^{i}_{t}$ are linearly transformed by $\tilde{{\bm{S}}}_{e}$. Please note that for each vector ${\bm{r}}^{i}_{t}$ (where $i=1,\dots,p$), its $j$-th element is scaled by the $j$-th column of the matrix $\tilde{{\bm{S}}}_{e}$. And the $j$-th elements of these vectors ${\bm{r}}^{i}_{t}$ all belong to the group ${\bm{X}}^{j}$. Therefore, our inductive gradient adjustments are generalized from one sampled data point of ${\bm{X}}^{j}$ to the remaining $(p-1)$ data points of ${\bm{X}}^{j}$, for all $j=1,\dots,n$.
3.3 Theoretical Analysis
In this subsection, we provide theoretical analysis to justify the feasibility of our IGA. In Theorem 3.1, we prove that adjustments based on eNTK matrix $\tilde{{\bm{K}}}$ asymptotically converge to those based on NTK matrix ${\bm{K}}$ as the network width increases. Then we delve into the estimate of training dynamics via inductive generalization of linear dynamics model of sampled data in Theorem 3.2.
Figure 2: Progressively amplified impacts on spectral bias of ReLU and SIREN by increasing the number of balanced eigenvalues via $\tilde{{\bm{S}}}_{e}$ when $p=8$, i.e., IGA. ReLU denotes the MLP with ReLU optimized via vanilla gradients; ReLU-$\tilde{{\bm{S}}}_{e}$-2 denotes the ReLU optimized via gradients adjusted by $\tilde{{\bm{S}}}_{e}$ with $2$ balanced eigenvalues. Details and more results are in the experiment 2 of Sec. 4 and Appendix B.
Figure 3: Relative error $\Delta_{k}$ curves of ReLU at 20Hz and SIREN at 40Hz with varying balanced eigenvalues and group size $p$. IGA-1 denotes that $p=1$, i.e., $\tilde{{\bm{K}}}$-based gradient adjustment; Baseline denotes vanilla gradients. The shaded area of lines indicates training fluctuations, represented by twice the standard deviation. Details and more results are in the experiment 2 of Sec. 4 and Appendix B.
In specific, we build upon previous training dynamics works (Du et al., 2018; Arora et al., 2019; Lee et al., 2019; Geifman et al., 2023) and analyze a two-layer network $f({\bm{x}};\Theta)$ with width $m$. Without loss of generality, we assume the training set $\{{{\bm{x}}_{i}},y_{i}\}_{i=1}^{N}$ and adopt the notations in Eq. 2-4. Detailed analysis settings are in Appendix A. Based on these settings, the following Theorem 3.1 show that $\tilde{{\bm{K}}}$-based adjustments achieves similar effects to ${\bm{K}}$-based adjustments.
Theorem 3.1.
(informal) Assuming that ${\bm{v}}_{i}^{\top}\tilde{{\bm{v}}}_{i}>0$ for $i\in[N]$. Let $\{g_{i}(x)\}_{i=1}^{N}$ be a set of Lipschitz continuous functions, with learning rate $\eta<\min\{(\max(g_{i}(\lambda))+\min(g_{i}(\lambda)))^{-1},(\max(g_{i}(\tilde{\lambda}))+\min(g_{i}(\tilde{\lambda})))^{-1}\}$, for all $\epsilon>0$, there always exists $M>0$, such that $m>M$, for $i\in[N]$, we have: $|g_{i}(\lambda_{i})-g_{i}(\tilde{\lambda_{i}})|<\epsilon_{1},\|{\bm{v}}_{i}-\tilde{{\bm{v}}}_{i}\|<\epsilon_{2}$, thus
$|(1-\eta g_{i}(\tilde{\lambda_{i}}))^{2}(\tilde{{\bm{v}}_{i}}^{\top}{\bm{r}}_{t})^{2}-(1-\eta g_{i}(\lambda_{i}))^{2}({\bm{v}}_{i}^{\top}{\bm{r}}_{t})^{2}|<\epsilon_{3},$
where $\epsilon_{1},\epsilon_{2},\epsilon_{3}$ and $\epsilon$ are of the same order as $\epsilon\rightarrow 0$.
Theorem 3.1 illustrates that eigenvalues and eigenvectors of $\tilde{{\bm{K}}}$ converge one-to-one to eigenvalues and eigenvectors of ${\bm{K}}$. That is, $\tilde{{\bm{S}}}$ impacts not only the convergence rates $\{g_{i}(\tilde{\lambda}_{i})\}_{i=1}^{N}$ similar to those of $S$ but also the corresponding convergence directions ($\tilde{{\bm{v}}}_{i}$). This theoretical results justify that the $\tilde{{\bm{K}}}$-based adjustment is approximately equivalent to ${\bm{K}}$-based adjustment. We further provide the empirical evidence in Sec. 4 and detailed proof of formal version in our Appendix A. Next, we analyze how training dynamics of ${\bm{X}}$ can be estimated from sampled data ${\bm{X}}_{e}$.
Theorem 3.2.
(informal) ${\bm{X}}$ is partitioned into $n$ groups ${\bm{X}}_{j}=\{{\bm{x}}^{j}_{1},\dots,{\bm{x}}^{j}_{p}\}$ and $N=np$. For ${\bm{X}}_{j}$, the sampled point denoted as ${\bm{x}}^{j}$ and these $n$ points form ${\bm{X}}_{e}=\{{\bm{x}}^{j}\}_{j=1}^{n},n\ll N$. There exists $\epsilon>0$, for $j\in[n]$ and $i_{1},i_{2}\in[p]$, such that $\max\{|(f({\bm{x}}^{j}_{i_{1}},\Theta_{t})-y^{j}_{i_{1}})-(f({\bm{x}}^{j}_{i_{2}},\Theta_{t})-y^{j}_{i_{2}})|,\|\nabla_{\Theta_{t}}f({\bm{x}}^{j}_{i_{1}})-\nabla_{\Theta_{t}}f({\bm{x}}^{j}_{i_{2}})\|\}<\epsilon$. Then, for any ${\bm{x}}_{i}$, assumed that ${\bm{x}}_{i}$ belongs to ${\bm{X}}_{j_{i}}$, such that:
$|(\Delta{\bm{r}}_{t})_{i}-p\nabla_{\Theta_{t}}f({\bm{x}}^{j_{i}})^{\top}\nabla_{\Theta_{t}}f({\bm{X}}_{e};\Theta_{t}){\bm{r}}^{t}_{e}|<\epsilon_{1}+\epsilon_{2},$
where $\epsilon_{1}=\Theta(\epsilon\frac{n^{3/2}}{\sqrt{m}})$; $\epsilon_{2}=\Theta(m^{-3/2})$; $\Theta(\cdot)$ denotes equivalence; $(\Delta{\bm{r}}_{t})_{i}=({\bm{r}}_{t+1}-{\bm{r}}_{t})_{i}$ denotes the residual of ${\bm{x}}_{i}$ at time step $t$; ${\bm{r}}_{e}^{t}$ denotes the residual on ${\bm{X}}_{e}$ at time step $t$; $\nabla_{\Theta_{t}}f({\bm{x}}^{j_{i}})^{\top}\nabla_{\Theta_{t}}f({\bm{X}}_{e};\Theta_{t})$ is the $j_{i}$-th row in $\tilde{{\bm{K}}}_{e}$.
Theorem 3.2 demonstrates that generalizing $\tilde{{\bm{K}}}_{e}$, induced from sampled data ${\bm{X}}_{e}$, to population data ${\bm{X}}$ provides an inductive estimation of overall training dynamics. The inductive estimate error $\epsilon_{1}+\epsilon_{2}$ decreases as the network width $m$ increases, the consistency bound $\epsilon$ decreases, or both. This theoretical insight encourages us to generalize the transformation matrix $\tilde{{\bm{S}}}_{e}$ to gradients of population data ${\bm{X}}$, thereby tailoring impacts on spectral bias via ${\bm{K}}_{e},\tilde{{\bm{S}}}_{e}$. We provide empirical validation of Theorem 3.2 in Sec. 4, with its formal version and proof in Appendix A.
@@ 3.4 Implementation Details
Sampling Strategy. As illustrated in Sec. 3.2, we need to sample ${\bm{X}}_{e}$ from groups of the population data ${\bm{X}}$ to obtain $\tilde{{\bm{K}}}_{e},\tilde{{\bm{S}}}_{e}$. In practice, we first form $n$ groups by simply considering $p$ adjacent input coordinates for each group. Specifically, for 1D and 2D inputs, groups are defined by non-overlapping intervals and patches, respectively. For $N$ $d$-dimensional input coordinates ($d\geq 3$), typically represented as a tensor of shape $(n_{1},\dots,n_{d},d)$, we flatten the first $d$ dimensions into a 2D tensor ${\bm{X}}^{\prime}$ of shape $(N,d)$ and then group it along the first dimension of ${\bm{X}}^{\prime}$ as 1D signals. Then, the input with the largest residuals in each group are selected to form the sampled data ${\bm{X}}_{e}$. To better illustrate the above process, we provide diagrams in Fig. 6. Details of hyperparameters $n$ and $p$ will be introduced in the experiment section. The ablation study of $n$ and $p$ is in the Appendix F. As shown in our wide experiments, the above strategy is simple yet effective and easy to implement.
@@
Construction Strategy. As illustrated in Eq. 4 and Theorem 3.2, we adjust the evenness of the spectrum of $\tilde{{\bm{K}}}_{e}$ via the transformation matrix $\tilde{{\bm{S}}}_{e}$ to purposefully improve the spectral bias. We have $\tilde{{\bm{K}}}_{e}=\sum_{i=1}^{n}\tilde{\tilde{\lambda}}_{i}\tilde{\tilde{{\bm{v}}_{i}}}\tilde{\tilde{{\bm{v}}_{i}}}^{\top}$ and assume that $\tilde{\tilde{\lambda}}_{1}>\dots>\tilde{\tilde{\lambda}}_{n}>0$. To adjust the evenness of spectrum of $\tilde{{\bm{K}}}_{e}$, $\tilde{{\bm{S}}}_{e}$ is constructed to balance the eigenvalues of different eigenvectors of $\tilde{{\bm{K}}}_{e}$; impacts are tailored by managing the number of balanced eigenvalues. For vanilla gradient descent, $\tilde{{\bm{S}}}_{e}$ is constructed as follows:
$\tilde{{\bm{S}}_{e}}=\sum_{i=start}^{end}(\tilde{\tilde{\lambda}}_{start}/\tilde{\tilde{\lambda}}_{i})\tilde{\tilde{\mathbf{v}}}_{i}\tilde{\tilde{\mathbf{v}}}_{i}^{\top}+\sum_{i\notin[start,end]}1\cdot\tilde{\tilde{\mathbf{v}}}_{i}\tilde{\tilde{\mathbf{v}}}_{i}^{\top}.$ (5)
Generally, $start$ is fixed at $1$; $end$ represents the controlled spectral range. Larger $end$ values have stronger impacts on spectral bias. We demonstrate this effect in experiment 2 of Sec. 4. For Adam, we utilize $\tilde{\tilde{\lambda}}_{end+1}$ instead of $\tilde{\tilde{\lambda}}_{start}$ to balance eigenvalues in Eq. 5 for better convergence, resulting in uniform eigenvalues of the modified $\tilde{{\bm{K}}}_{e}$ across the $(end-start+2)$ eigenvectors. Then, momentum and adaptive learning rates are updated according to the adjusted gradients. More discussion are in Appendix D.
@@
Multidimensional Output Approximation. In practical INRs tasks, the output is multidimensional, such as for color images. Compared to the data size $N$, it is trivial but still affects efficiency. Thus, we simply compute the Jacobian matrix $\nabla_{\Theta}f({\bm{X}};\Theta)$ by sum-of-logits, which is shown to be efficient in Mohamadi et al. (2023); Shi et al. (2024b).
Figure 4: Visual examples of 2D color image approximation results by different training dynamics methods. Enlarged views of regions labeled by red boxes are provided. The residuals of these regions in the Fourier domain are visualized through heatmaps in the top right corner. The increase in error corresponds to the transition of colors in the heatmaps from blue to red. Detailed settings are in Sec. 5.1.
Table 1: Average metrics of 2D color image approximation results by different methods. The detailed settings can be found in Sec. 5.1. Per-scene results are provided in Table 8-11. Note: SIREN+BN is omitted due to our finding of their incompatibility, neither does the original work (Cai et al., 2024) demonstrate the feasibility.
Average ReLU PE SIREN
Metric Vanilla +FR +BN +IGA Vanilla +FR +BN +IGA Vanilla +FR +IGA
PSNR $\uparrow$ 21.78 22.14 22.55 23.00 28.64 31.65 28.78 32.46 32.65 32.61 33.48
SSIM $\uparrow$ 0.4833 0.4919 0.5004 0.5126 0.7832 0.8167 0.8030 0.8822 0.8975 0.8991 0.9121
MS-SSIM $\uparrow$ 0.6521 0.6800 0.7090 0.7383 0.9466 0.9564 0.9554 0.9752 0.9818 0.9820 0.9847
LPIPS $\downarrow$ 0.6302 0.6315 0.6182 0.5549 0.2223 0.1869 0.2346 0.0938 0.0807 0.0813 0.0668